# Safety Mirage: How Spurious Correlations Undermine VLM Safety Fine-Tuning and Can Be Mitigated by Machine Unlearning

### v2

@@ Abstract
Recent vision language models (VLMs) have made remarkable strides in generative modeling with multimodal inputs, particularly text and images. However, their susceptibility to generating harmful content when exposed to unsafe queries raises critical safety concerns. While current alignment strategies primarily rely on supervised safety fine-tuning with curated datasets, we identify a fundamental limitation we call the “safety mirage”, where supervised fine-tuning inadvertently reinforces spurious correlations between superficial textual patterns and safety responses, rather than fostering deep, intrinsic mitigation of harm. We show that these spurious correlations leave fine-tuned VLMs vulnerable even to a simple one-word modification-based attack, where substituting a single word in text queries with a spurious correlation-inducing alternative can effectively bypass safeguards. Additionally, these correlations contribute to the over-prudence, causing fine-tuned VLMs to refuse benign queries unnecessarily. To address these issues, we show machine unlearning (MU) as a powerful alternative to supervised safety fine-tuning, as it avoids biased feature-label mappings and directly removes harmful knowledge from VLMs while preserving their general capabilities. Extensive evaluations across safety benchmarks show that under MU-based alignment reduces the attack success rate by up to 60.17% and cuts unnecessary rejections by over 84.20%. Codes are available at https://github.com/OPTML-Group/VLM-Safety-Unlearn. WARNING: There exist AI generations that may be offensive in nature.
@@ 1 Introduction
Recent multi-modal models have achieved advancements by integrating text and images (Alayrac et al., 2022; Awadalla et al., 2023; Hurst et al., 2024; Gao et al., 2023; Li et al., 2023a). A prevalent architecture in VLMs maps visual embeddings into the language model’s latent space via a dedicated projection module (Liu et al., 2023, 2024a; Zhu et al., 2023; Chen et al., 2023a; Wang et al., 2024a). While the large language model (LLM) backbone effectively integrates projected visual information, the image modality also introduces new vulnerabilities (Guo et al., 2024; Qi et al., 2023a; Liu et al., 2024b, c; Gong et al., 2023). Specifically, images can function as a “foreign language” (Pi et al., 2024), creating pathways for unsafe input queries, even when the underlying LLM has been aligned for safety (Pi et al., 2024; Chakraborty et al., 2024; Ding et al., 2025). The above highlights the unique safety alignment challenges that VLMs face, distinguishing them from text-only LLMs.
Despite the emergence of safety challenges in VLMs, recent studies have revealed a surprising empirical finding: Enhancing VLM safety could be as simple as applying supervised fine-tuning (SFT), assuming access to a high-quality, dual-modality curated safety fine-tuning dataset (Liu et al., 2024c; Zhou et al., 2024; Luo et al., 2024; Zhang et al., 2024a; Gu et al., 2024). One compelling piece of evidence is that fine-tuning on widely used VLM safety datasets such as VLGuard (Zong et al., 2024) and SPA-VL (Zhang et al., 2024a), which substantially improves robustness against unsafe queries and jailbreaking attacks. The surprising effectiveness of safety SFT has sparked growing interest in re-evaluating its reliability. However, a key limitation of such safety fine-tuning is the issue of over-prudence (Ding et al., 2025; Guo et al., 2024; Ding et al., 2024), where VLMs become overly conservative and unnecessarily reject benign queries.
@@
Figure 1: Schematic overview of safety mirage findings of safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed fine-tuned on VLGuard (Zong et al., 2024)) (a) One-word attack vulnerability: A minor modification (e.g., replacing the first instruction word “Share” with “What” in an unsafe query) bypasses the safety mechanism learned from fine-tuning on VLGuard, even though the model correctly rejects the original query. (b) Over-prudence issue: Similar to the one-word attack, a minor modification by replacing “What” with “Share” can cause unnecessary refusals even for benign queries. (c) Root cause of spurious correlations from fine-tuning dataset biases: Certain words are disproportionately linked with specific safety labels, such as “Share” strongly correlated with rejection, while “What” highly associated with non-rejection. (d) Effectiveness of unlearning-based safety fine-tuning: Using unlearning methods NPO (Zhang et al., 2024b) and RMU (Li et al., 2024a) to do the safety fine-tune on the VLM, which could enhance robustness against attacks and reduce over-prudence, outperforming both the original model LLaVA-v1.5-7B (“Original”) and the supervised fine-tuned LLaVA-v1.5-7B-Mixed (Zong et al., 2024) (“Prior work”).
@@
Building on our identification of spurious correlations in VLMs safety fine-tuning, we propose improving current safety fine-tuning approaches through machine unlearning (MU). Originally designed to remove the influence of undesired data or knowledge from ML models, MU preserves essential knowledge while avoiding unintended disruptions to causally unrelated information (Liu et al., 2025; Cao and Yang, 2015; Bourtoule et al., 2021). We propose adapting MU to VLMs safety fine-tuning as a more robust alternative to traditional supervised approaches. Instead of enforcing safety through direct supervision, MU enhances VLM safety by erasing the influence of unsafe knowledge in a label-free manner, thereby preventing the formation of spurious correlations between input features and safety labels. Although MU has been applied to VLMs safety fine-tuning in prior work (Chakraborty et al., 2024; Chen et al., 2025; Huo et al., 2025), its unique advantage in mitigating spurious correlations within fine-tuning datasets remains unexplored. Fig. 1(d) illustrates the effectiveness of adapting two LLM unlearning approaches, NPO (negative preference optimization) (Zhang et al., 2024b) and RMU (representation misdirection for unlearning) (Li et al., 2024a), in enhancing robustness against jailbreaking attacks and reducing over-prudence rates.
@@
④We conduct extensive experiments across multiple VLMs safety evaluation benchmarks, including VLGuard (Zong et al., 2024), SPA-VL (Zhang et al., 2024a), MM-SafetyBench (Liu et al., 2024c), and FigStep (Gong et al., 2023), and assess model utility on standard VQA datasets. Our results confirm the safety mirage phenomenon and demonstrate that MU-based safety fine-tuning effectively mitigates spurious correlations and reduces over-prudence.
@@ 2 Related Work
VLM safety: Attack and defense. With the rapid advancement of VLMs (Liu et al., 2023, 2024a; Zhu et al., 2023; Ye et al., 2023; Wang et al., 2023; Li et al., 2023b; Alayrac et al., 2022; Awadalla et al., 2023; Gao et al., 2023), safety concerns have become increasingly prominent due to their potential to generate harmful or inappropriate content. While LLMs have been extensively studied for safety risks, leading to the development of attack strategies (Yang et al., 2023; Wei et al., 2023a; Huang et al., 2023; Shu et al., 2023), defense mechanisms (Li et al., 2023c; Cao et al., 2023; Kumar et al., 2023), and robust evaluation datasets (Bianchi et al., 2023; Li et al., 2024b; Ji et al., 2023), VLMs introduce additional challenges due to the complexity of multimodal inputs (Pi et al., 2024; Chakraborty et al., 2024; Ding et al., 2025), making them even more vulnerable to jailbreaking and adversarial manipulation (Guo et al., 2024; Qi et al., 2023a; Liu et al., 2024b, c; Gong et al., 2023). Attacks on VLMs often leverage the dual-modality nature of these models. One approach embeds unsafe textual queries into images through typographic manipulation, enabling the model to bypass safety filters and generate harmful outputs (Gong et al., 2023; Liu et al., 2024c). Another strategy involves using gradient-based adversarial image generation (Bailey et al., 2023; Dong et al., 2023; Luo et al., 2023; Qi et al., 2023b; Zhao et al., 2023) to trigger harmful responses, demonstrating that VLMs remain susceptible to adversarial perturbations despite safety fine-tuning. Defensive strategies for VLM safety generally fall into two categories: Inference-time defenses and fine-tuning with curated safety datasets. The former aligns safety responses dynamically at runtime, mitigating unsafe outputs using various filtering and rejection mechanisms (Wang et al., 2024b; Chen et al., 2023b; Pi et al., 2024; Gou et al., 2024; Ding et al., 2024). The latter focuses on red-teaming dataset curation (Liu et al., 2024c; Zhou et al., 2024; Luo et al., 2024; Zhang et al., 2024a; Gu et al., 2024; Zong et al., 2024; Li et al., 2024c), enabling VLMs to be explicitly trained to reject harmful content while retaining utility for benign tasks.
@@
Machine unlearning in VLMs. MU (Liu et al., 2025; Cao and Yang, 2015; Bourtoule et al., 2021) is designed to remove harmful data influences from a pre-trained model while preserving its overall utility. In the LLM domain, recent work has explored targeted forgetting techniques to erase specific knowledge without compromising performance (Zhang et al., 2024b; Li et al., 2024a; Yao et al., 2024). In VLMs, several benchmarks have established systematic evaluation frameworks for MU algorithms (Liu et al., 2024d; Dontsov et al., 2024; Ma et al., 2024). For safety fine-tuning, prior studies have applied MU-based approaches to mitigate harmful content generation (Chen et al., 2025; Huo et al., 2025; Chakraborty et al., 2024). Our work builds on these efforts by leveraging MU to specifically mitigate spurious correlations present in safety fine-tuning datasets.
@@ 3 Preliminary and Problem Statement
Previous works (Pi et al., 2024; Gong et al., 2023; Liu et al., 2024c) highlight the importance of safety alignment in VLMs across both textual and visual modalities. Consequently, many efforts have focused on curating high-quality dual-modality safety datasets for VLMs. For instance, VLGuard (Zong et al., 2024) covers various text-image scenarios, including unsafe cases where the text is unsafe or both modalities are unsafe, as well as safe cases where both are benign. SPA-VL (Zhang et al., 2024a) provides large-scale preference data in the form of question–response tuples, where each image sample pairs a safety-related question with a chosen response and a rejected response. Leveraging these curated safety datasets, recent works (Chakraborty et al., 2024; Ding et al., 2025) show that simple fine-tuning approaches on such datasets can yield surprisingly strong safety performance, even against common jailbreaking attacks (Zou et al., 2023; Wei et al., 2023b; Röttger et al., 2023). In this work, we revisit the VLM safety fine-tuning problem and later argue that the observed safety improvements from fine-tuning may be an illusion.
@@
Figure 2: Visualization of question-answer samples on the safety fine-tuned VLMs (LLaVA-v1.5-7B-Mixed and LLaVA-v1.5-7B-Posthoc (Zong et al., 2024; Taori et al., 2023)). Green shield represents the correct response, either a safe rejection for harmful queries or a valid answer for benign queries. Red exclamation ! indicates an unsafe response to harmful queries. And red question ? represents an inappropriate rejection for a safe query. (a) Successfully jailbreak: The safety fine-tuned model originally produces rejection-based responses for unsafe queries. However, asking the initial question by starting with the word with “What” can easily bypass this safeguard. (b) Over-prudence: A minor modification by asking the problem start from “What” with “Share” can trigger refusals even for benign queries.
Safety mirage: Motivation and problem of interest. Although safety fine-tuned VLMs could exhibit over-rejection, refusing even benign queries (Ding et al., 2025; Guo et al., 2024; Ding et al., 2024), their safety against unsafe queries remains highly robust under common jailbreaking attacks (Zong et al., 2024; Zhang et al., 2024a). The seemingly ‘robust’ safety observed after fine-tuning motivates us to re-examine its true reliability. As demonstrated in Fig. 1, the current safety fine-tuned VLM remains highly vulnerable to simple paraphrasing of text queries even when only the first question word is modified. As a supplement, Fig. 2 & A1 provides motivating examples illustrating the vulnerability of safety fine-tuned VLMs to both jailbreaking attacks and over-prudence. For example, for the VLGuard fine-tuned models, unsafe queries prefixed with the innocuous question word “What” successfully bypass the safeguard, and harmless prompts start with “Share” trigger over-rejection effect. Notably, the choice of the question word is not random but rather stems from the spurious correlations embedded in the safety fine-tuning dataset, as we will illustrate in Sec. 4.
@@ 4 The Risks of Spurious Correlations in VLM Safety Fine-Tuning
Refer to caption Refer to caption Refer to caption
At first glance, the use of safety labels appears appropriate. However, hidden bias may arise when safety labels become strongly correlated with spurious features in the input data, particularly within textual queries, which are the focus of this work. Here the term “spurious features” refers to non-essential features in inputs (primarily for texts in this work) that do not contribute to the fundamental meaning or task-relevance of the input query, in contrast to the “core” features. For example, in Fig. 1, the initial words “What” or “Share” function as spurious features, since they are not directly related to the query’s actual content and can be easily substituted with other question words. In contrast, core features (such as “crime”) are more informative, representing content-related words that capture the true meaning of the query. Therefore, we define spurious correlations as the (unexpected) strong associations between spurious input features and the safety labels in the fine-tuning dataset. The above conceptualization of spurious correlations is inspired by and remains consistent with conventional spurious correlation analyses in image classification (Sagawa et al., 2020), where background pixels serve as spurious features while object pixels represent core content.
@@
Figure 5: $R$ vs. $K$-shot one-word modification, evaluated before and after applying the “Share” initialized one-word modification to safe input queries, causes the over-prudence phenomenon of the safety fine-tuned VLM LLaVA-v1.5-7B-Mixed model (Zong et al., 2024) on VLGuard safe.

### v3

# arXiv:2503.11832v3[cs.AI] 02 Mar 2026

@@ Abstract
Recent vision language models (VLMs) have made remarkable strides in generative modeling with multimodal inputs, particularly text and images. However, their susceptibility to generating harmful content when exposed to unsafe queries raises critical safety concerns. While current alignment strategies primarily rely on supervised safety fine-tuning with curated datasets, we identify a fundamental limitation we call the “safety mirage,” where supervised fine-tuning inadvertently reinforces spurious correlations between superficial textual patterns and safety responses, rather than fostering deep, intrinsic mitigation of harm. We show that these spurious correlations leave fine-tuned VLMs vulnerable even to a simple one-word modification-based attack, where substituting a single word in text queries with a spurious correlation-inducing alternative can effectively bypass safeguards. Additionally, these correlations contribute to over-prudence, causing fine-tuned VLMs to refuse benign queries unnecessarily. To address these issues, we show machine unlearning (MU) as a powerful alternative to supervised safety fine-tuning, as it avoids biased feature-label mappings and directly removes harmful knowledge from VLMs while preserving their general capabilities. Extensive evaluations across safety benchmarks show that under MU-based alignment reduces the attack success rate by up to 60.27% and cuts unnecessary rejections by over 84.20%. Codes are available at https://github.com/OPTML-Group/VLM-Safety-Unlearn. WARNING: There exist AI generations that may be offensive in nature.
@@ 1 Introduction
Recent multi-modal models have achieved advancements by integrating text and images (Alayrac et al., 2022; Awadalla et al., 2023; Hurst et al., 2024; Gao et al., 2023; Li et al., 2023b). A prevalent architecture in VLMs maps visual embeddings into the language model’s latent space via a dedicated projection module (Liu et al., 2023; 2024a; Zhu et al., 2023; Chen et al., 2023a; Wang et al., 2024a). While the large language model (LLM) backbone effectively integrates projected visual information, the image modality also introduces new vulnerabilities (Guo et al., 2024; Qi et al., 2023b; Liu et al., 2024c; b; Gong et al., 2023). Specifically, images can function as a “foreign language” (Pi et al., 2024), creating pathways for unsafe input queries, even when the underlying LLM has been aligned for safety (Pi et al., 2024; Chakraborty et al., 2024; Ding et al., 2025). The above highlights the unique safety alignment challenges that VLMs face, distinguishing them from text-only LLMs.
Despite the emergence of safety challenges in VLMs, recent studies have revealed a surprising empirical finding: Enhancing VLM safety could be as simple as applying supervised fine-tuning (SFT), assuming access to a high-quality, dual-modality curated safety fine-tuning dataset (Liu et al., 2024b; Zhou et al., 2024; Luo et al., 2024; Zhang et al., 2024b; Gu et al., 2024). One compelling piece of evidence is that fine-tuning on widely used VLM safety datasets such as VLGuard (Zong et al., 2024) and SPA-VL (Zhang et al., 2024b), which substantially improves robustness against unsafe queries and jailbreaking attacks. The surprising effectiveness of safety SFT has sparked growing interest in re-evaluating its reliability. However, a key limitation of such safety fine-tuning is the issue of over-prudence (Ding et al., 2025; Guo et al., 2024; Ding et al., 2024), where VLMs become overly conservative and unnecessarily reject benign queries.
@@
Figure 1: Schematic overview of safety mirage findings of safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed fine-tuned on VLGuard (Zong et al., 2024)) (a) One-word attack vulnerability: A minor modification (e.g., replacing the first instruction word “Share” with “What” in an unsafe query) bypasses the safety mechanism learned from fine-tuning on VLGuard, even though the model correctly rejects the original query. (b) Over-prudence issue: Similar to the one-word attack, a minor modification by replacing “What” with “Share” can cause unnecessary refusals even for benign queries. (c) Root cause of spurious correlations from fine-tuning dataset biases: Certain words are disproportionately linked with specific safety labels, such as “Share” strongly correlated with rejection, while “What” highly associated with non-rejection. (d) Effectiveness of unlearning-based safety fine-tuning: The unlearning methods NPO (Zhang et al., 2024a) and RMU (Li et al., 2024c), enhance robustness against attacks and reduce over-prudence, outperforming both the original model LLaVA-v1.5-7B (“Original”) and the supervised fine-tuned LLaVA-v1.5-7B-Mixed (Zong et al., 2024) (“Prior work”).
@@
Building on our identification of spurious correlations in VLMs safety fine-tuning, we propose improving current safety fine-tuning approaches through machine unlearning (MU). Originally designed to remove the influence of undesired data or knowledge from ML models, MU preserves essential knowledge while avoiding unintended disruptions to causally unrelated information (Liu et al., 2025; Cao and Yang, 2015; Bourtoule et al., 2021). We propose adapting MU to VLMs safety fine-tuning as a more robust alternative to traditional supervised approaches. Instead of enforcing safety through direct supervision, MU enhances VLM safety by erasing the influence of unsafe knowledge in a label-free manner, thereby preventing the formation of spurious correlations between input features and safety labels. Although MU has been applied to VLMs safety fine-tuning in prior work (Chakraborty et al., 2024; Chen et al., 2025; Huo et al., 2025), its unique advantage in mitigating spurious correlations within fine-tuning datasets remains unexplored. Fig. 1(d) illustrates the effectiveness of adapting two LLM unlearning approaches, NPO (negative preference optimization) (Zhang et al., 2024a) and RMU (representation misdirection for unlearning) (Li et al., 2024c), in enhancing robustness against jailbreaking attacks and reducing over-prudence rates.
@@
④We conduct extensive experiments across multiple VLMs safety evaluation benchmarks, including VLGuard (Zong et al., 2024), SPA-VL (Zhang et al., 2024b), MM-SafetyBench (Liu et al., 2024b), and FigStep (Gong et al., 2023), and assess model utility on standard VQA datasets. Our results confirm the safety mirage phenomenon and demonstrate that MU-based safety fine-tuning effectively mitigates spurious correlations and reduces over-prudence.
@@ 2 Related Work
VLM safety: Attack and defense. With the rapid advancement of VLMs (Liu et al., 2023; 2024a; Zhu et al., 2023; Ye et al., 2023; Wang et al., 2023; Li et al., 2023a; Alayrac et al., 2022; Awadalla et al., 2023; Gao et al., 2023), safety concerns have become increasingly prominent due to their potential to generate harmful or inappropriate content. While LLMs have been extensively studied for safety risks, leading to the development of attack strategies (Yang et al., 2023; Wei et al., 2023b; Huang et al., 2023; Shu et al., 2023), defense mechanisms (Li et al., 2023c; Cao et al., 2023; Kumar et al., 2023), and robust evaluation datasets (Bianchi et al., 2023; Li et al., 2024a; Ji et al., 2023), VLMs introduce additional challenges due to the complexity of multimodal inputs (Pi et al., 2024; Chakraborty et al., 2024; Ding et al., 2025), making them even more vulnerable to jailbreaking and adversarial manipulation (Guo et al., 2024; Qi et al., 2023b; Liu et al., 2024c; b; Gong et al., 2023). Attacks on VLMs often leverage the dual-modality nature of these models. One approach embeds unsafe textual queries into images through typographic manipulation, enabling the model to bypass safety filters and generate harmful outputs (Gong et al., 2023; Liu et al., 2024b). Another strategy involves using gradient-based adversarial image generation (Bailey et al., 2023; Dong et al., 2023; Luo et al., 2023; Qi et al., 2023a; Zhao et al., 2023) to trigger harmful responses, demonstrating that VLMs remain susceptible to adversarial perturbations despite safety fine-tuning. Defensive strategies for VLM safety generally fall into two categories: Inference-time defenses and fine-tuning with curated safety datasets. The former aligns safety responses dynamically at runtime, mitigating unsafe outputs using various filtering and rejection mechanisms (Wang et al., 2024b; Chen et al., 2023b; Pi et al., 2024; Gou et al., 2024; Ding et al., 2024). The latter focuses on red-teaming dataset curation (Liu et al., 2024b; Zhou et al., 2024; Luo et al., 2024; Zhang et al., 2024b; Gu et al., 2024; Zong et al., 2024; Li et al., 2024b), enabling VLMs to be explicitly trained to reject harmful content while retaining utility for benign tasks.
@@
Machine unlearning in VLMs. MU (Liu et al., 2025; Cao and Yang, 2015; Bourtoule et al., 2021) is designed to remove harmful data influences from a pre-trained model while preserving its overall utility. In the LLM domain, recent work has explored targeted forgetting techniques to erase specific knowledge without compromising performance (Zhang et al., 2024a; Li et al., 2024c; Yao et al., 2024). In VLMs, several benchmarks have established systematic evaluation frameworks for MU algorithms (Liu et al., 2024d; Dontsov et al., 2024; Ma et al., 2024). For safety fine-tuning, prior studies have applied MU-based approaches to mitigate harmful content generation (Chen et al., 2025; Huo et al., 2025; Chakraborty et al., 2024). Our work builds on these efforts by leveraging MU to specifically mitigate spurious correlations present in safety fine-tuning datasets.
Figure 2: Visualization of question-answer samples on the safety fine-tuned VLMs (LLaVA-v1.5-7B-Mixed and LLaVA-v1.5-7B-Posthoc (Zong et al., 2024; Taori et al., 2023)). Green shield represents the correct response, either a safe rejection for harmful queries or a valid answer for benign queries. Red exclamation ! indicates an unsafe response to harmful queries. And red question ? represents an inappropriate rejection for a safe query. (a) Successfully jailbreak: The safety fine-tuned model originally produces rejection-based responses for unsafe queries. However, asking the initial question by starting with the word with “What” can easily bypass this safeguard. (b) Over-prudence: A minor modification by asking the problem start from “What” with “Share” can trigger unnecessary refusals even for benign queries.
@@ 3 Preliminary and Problem Statement
Previous works (Pi et al., 2024; Gong et al., 2023; Liu et al., 2024b) highlight the importance of safety alignment in VLMs across both textual and visual modalities. Consequently, many efforts have focused on curating high-quality dual-modality safety datasets for VLMs. For instance, VLGuard (Zong et al., 2024) covers various text-image scenarios, including unsafe cases where the text is unsafe or both modalities are unsafe, as well as safe cases where both are benign. SPA-VL (Zhang et al., 2024b) provides large-scale preference data in the form of question–response tuples, where each image sample pairs a safety-related question with a chosen response and a rejected response. Leveraging these curated safety datasets, recent works (Chakraborty et al., 2024; Ding et al., 2025) show that simple fine-tuning approaches on such datasets can yield surprisingly strong safety performance, even against common jailbreaking attacks (Zou et al., 2023; Wei et al., 2023a; Röttger et al., 2023). In this work, we revisit the VLM safety fine-tuning problem and later argue that the observed safety improvements from fine-tuning may be an illusion.
@@
Refer to caption Refer to caption Refer to caption
Safety mirage: Motivation and problem of interest. Although safety fine-tuned VLMs could exhibit over-rejection, refusing even benign queries (Ding et al., 2025; Guo et al., 2024; Ding et al., 2024), their safety against unsafe queries remains highly robust under common jailbreaking attacks (Zong et al., 2024; Zhang et al., 2024b). The seemingly ‘robust’ safety observed after fine-tuning motivates us to re-examine its true reliability. As demonstrated in Fig. 1, the current safety fine-tuned VLM remains highly vulnerable to simple paraphrasing of text queries even when only the first question word is modified. As a supplement, Fig. 2 & A1 provides motivating examples illustrating the vulnerability of safety fine-tuned VLMs to both jailbreaking attacks and over-prudence. For example, for the VLGuard fine-tuned models, unsafe queries prefixed with the innocuous question word “What” successfully bypass the safeguard, and harmless prompts start with “Share” trigger over-rejection effect. Notably, the choice of the question word is not random but rather stems from the spurious correlations embedded in the safety fine-tuning dataset, as we will illustrate in Sec. 4.
@@ 4 The Risks of Spurious Correlations in VLM Safety Fine-Tuning
At first glance, the use of safety labels appears appropriate. However, hidden bias may arise when safety labels become strongly correlated with spurious features in the input data, particularly within textual queries, which are the focus of this work. Here the term “spurious features” refers to non-essential features in inputs (primarily for texts in this work) that do not contribute to the fundamental meaning or task-relevance of the input query, in contrast to the “core” features. For example, in Fig. 1, the initial words “What” or “Share” function as spurious features, since they are not directly related to the query’s actual content and can be easily substituted with other question words. In contrast, core features (such as “crime”) are more informative, representing content-related words that capture the true meaning of the query. Therefore, we define spurious correlations as the (unexpected) strong associations between spurious input features and the safety labels in the fine-tuning dataset. The above conceptualization of spurious correlations is inspired by conventional spurious correlation analyses in image classification (Sagawa et al., 2020), where background pixels serve as spurious features while object pixels represent core content. In addition, the introduction of spurious correlations during safety alignment echoes the alignment tax (Siddiqui et al., 2026) and emergent misalignment phenomena (Betley et al., 2026), wherein fine-tuning a broadly capable pretrained model on a narrow objective can inadvertently induce new, unintended misaligned behaviors.
@@
Figure 5: Rejection rate vs. $K$-shot one-word modification, evaluated before and after applying the “Share” initialized one-word modification to safe input queries, causes the over-prudence phenomenon of the safety fine-tuned VLM LLaVA-v1.5-7B-Mixed model (Zong et al., 2024) on VLGuard safe.